# Approximate Sequential Optimization for Informative Path Planning

### v1

@@ Abstract
We consider the problem of finding an informative path through a graph, given initial and terminal nodes and a given maximum path length. We assume that a linear noise corrupted measurement is taken at each node of an underlying unknown vector that we wish to estimate. The informativeness is measured by the reduction in uncertainty in our estimate, evaluated using several metrics. We present a convex relaxation for this informative path planning problem, which we can readily solve to obtain a bound on the possible performance. We develop an approximate sequential method where the path is constructed segment by segment through dynamic programming. This involves solving an orienteering problem, with the node reward acting as a surrogate for informativeness, taking the first step, and then repeating the process. The method scales to very large problem instances and achieves performance not too far from the bound produced by the convex relaxation. We also demonstrate our method’s ability to handle adaptive objectives, multimodal sensing, and multi-agent variations of the informative path planning problem.
@@ 1 Introduction
A variety of real-world problems can be formulated as IPP problems, such as planetary rover exploration, search and rescue, and environmental monitoring. IPP merges the fields of robotics, artificial intelligence, and spatial data analysis. It carries substantial implications for a myriad of applications, where efficient data collection from a potentially hazardous or inaccessible environment is essential. However, the problem presents significant challenges. The IPP problem is known to be NP-hard due to its combinatorial nature with a vast search space for potential solutions [3]. Successful solutions typically require reasoning about the entire trajectory; myopic approaches often result in inefficient paths, causing the agent to repeat exploration of the same areas without using the full resource budget [4].
@@
The IPP problem is closely related to the optimal design of experiments and sensor selection. [1] presented a convex relaxation of the sensor selection problem that involves choosing a set of $k$ sensor measurements, from a set of $m$ possible or potential sensor measurements, that minimizes the error in estimating some parameters [1]. The sensor selection problem is closely related to the IPP problem. However, sensor selection does not consider the movement constraints of the agent so the rounding approach presented by [1] is not directly applicable to the IPP problem. [5] also considered the sensor placement problem and presented several approximate solution methods [5]. [6] introduced an approximate algorithm for both the single and multi-agent IPP problems using Gaussian processes as the underlying environment model [6]. [7] used hybrid search to solve risk-bound goal-directed planning problems, focusing on complex 3D environments and the modeling of trajectory risk [7]. While [8] focused on approximability for cost-optimal planning [8].
@@
[2] introduced a mixed-integer formulation for the discrete IPP problem in random fields and aimed to optimize over the collected measurement subset and all linear estimators [2]. The approach presented by [2] can produce globally optimal solutions on relatively small graphs; however, the mixed integer program fails to scale to larger graph sizes which is often necessary in practice for planning at higher resolutions [2]. [9] used an evolutionary strategy to optimize a parameterized path in continuous space [9]. [10] formulated the problem as a mixed integer semidefinite program and used a receding horizon optimal control approach based on the information filter to produce a solution [10]. [11] used a receding horizon control strategy with a discrete set of motion primitives to evaluate the objective and then select the best candidate, execute the action, update the Gaussian process belief, and then replan [11].
@@
Another common approach is to formulate the IPP problem as a partially observable Markov decision process (POMDP). [12] formulated the IPP problem as a POMDP and solved it using sequential Bayesian optimization through Monte Carlo tree search with upper confidence bound for trees (MCTS-UCT) [12]. This work has been extended to modify the reward function for achieving monitoring behavior that exploits areas with high gradients, and for additional reasoning over continuous action spaces through Bayesian optimization [13, 14]. [15], on the other hand, used partially observable Monte Carlo planning (POMCP) with Gaussian process beliefs for estimating quantiles of the underlying world state [15]. Their methodology proposes sample locations for a team of scientists to investigate following their exploration completion. [16] presented a hybrid approach that combines tree search with an offline-learned neural network to predict informative sensing actions [16]. [17] focused on the multimodal sensing extension of the IPP problem where the agent has multiple sensors to choose from and solved it with tree search [17].
@@
These approaches all produce reasonable solutions on graphs with a relatively small number of nodes; however, they struggle to scale to large graphs with higher resolution which is necessary for many real-world applications. Additionally, MIPs are not well-suited to handle adaptive objectives where the objective function depends on the history of measurement values received. Our results demonstrate that our approach is able to scale to larger problem instances for the IPP problem, while also handling adaptive objectives such as the expected improvement metric, and accounting for multimodal sensing and multi-agent variations of the IPP problem [18, 17]. We directly compare our approach with those of [2] and [17] [2, 17].
@@ 2.1 Feasible Paths
We model the environment as a directed graph with vertices $V=\{1,\dots,n\}$, edges $\mathcal{E}\subseteq V\times V$. Our goal is to find a feasible path $p$ from the start vertex $s$ to the goal vertex $g$ where we assume that such a feasible path exists. We denote a path of length $L$ as $p=(i_{1},i_{2},\dots,i_{L})$ where $i_{1}=s$ and $i_{L}=g$. A feasible path is one that starts at $s$ and ends at $g$ while remaining within the given maximum path length $L^{\text{max}}$; that is, $L\leq L^{\text{max}}$.
@@ 2.2 Sensing Along a Path
$\hat{x}=\bar{x}+\left(\sigma^{-2}\sum_{j\in p}a_{j}a_{j}^{T}+\Sigma_{x}^{-1}\right)^{-1}\sigma^{-2}\sum_{k=1}^{L}(y_{k}-\bar{y}_{k})a_{i_{k}}$ (3)
@@
$\Sigma=\left(\sigma^{-2}\sum_{j\in p}a_{j}a_{j}^{T}+\Sigma_{x}^{-1}\right)^{-1}$ (4)
where $\bar{x}$ is the prior mean and $\bar{y}_{k}=a_{i_{k}}^{T}\bar{x}$. Note that this is exactly the Gaussian process mean and covariance [19].
@@ 2.3 Measures of Informativeness
Note that each scalar measure of informativeness is a function of the eigenvalues of $\Sigma$. We will denote the measure of informativeness of a path as $\phi(p)$ where smaller values of $\phi(p)$ correspond to more informative paths.
@@ 3 Mixed Integer Convex Formulation
$\Sigma_{z}=\left(\sigma^{-2}\sum_{i=1}^{n}\left(\sum_{j\in N_{out}(i)}z_{ij}\right)a_{i}a_{i}^{T}+\Sigma_{x}^{-1}\right)^{-1}.$ (15)
With some slight abuse of notation we can then write our parameterized objective as $\phi(z)$. We then have the following convex mixed integer problem:
@@
While it is not immediately obvious, the $\phi(z)$ objective is convex in both the A, B, and D-IPP objectives introduced in Section 2.3 as discussed in section 4.6.2 of [21]. The problem would be entirely convex if it were not for the integer constraints on $z$ in Equation 14.
@@ 5.2 Additional Objectives
$\mathbb{E}[I_{p_{i}}(y)]=(y_{\text{min}}-\hat{\mu}_{i})\Phi\left(\frac{y_{\text{min}}-\hat{\mu}_{i}}{\hat{\sigma}_{i}}\right)+\hat{\sigma}_{i}^{2}\mathcal{N}(y_{\text{min}}\mid\hat{\mu}_{i},\hat{\sigma}_{i}^{2})$
where
$I_{p_{i}}(y)=\begin{cases}y_{\min}-y&\text{ if }y<y_{\min}\\ 0&\text{ otherwise }\end{cases}$
is the improvement over the current minimum value $y_{\text{min}}$ found so far, $\Phi(\cdot)$ is the cumulative distribution function, and $\hat{\mu}_{i}=\hat{x}_{i}$ with $\hat{\sigma}_{i}^{2}=\Sigma_{ii}$ are the predicted mean and variance of the Gaussian process at node $i$ in Equation 3 and Equation 4 respectively [19]. Our problem then becomes
minimize $\sum_{i=1}^{m}\mathbb{E}[I_{p_{i}}(y)]$
@@
This is a non-convex objective and is not easily handled by current optimizers. However, our approximate sequential path optimization approach can handle this objective by computing $r_{j}=\mathbb{E}[I_{p_{j}}(y)]$ for each of the $m$ prediction vertices in the environment with the current values of $\hat{x}$ and $\Sigma$ obtained from the current path $p_{t}$. Note that we want to visit locations in the environment with high expected improvement to minimize the net expected improvement left in the environment so there is no negative sign in the expression for $r_{j}$
@@ 5.3 Multimodal Sensing
To solve the AIPPMS problem, we first solve the AIPP problem with our approximate sequential path optimization approach and the respective objective presented in Section 4. We then use the resulting path as the starting point for the multimodal sensor selection problem. The multimodal sensor selection problem is a variant of the sensor selection problem presented by [1]. In this case, we assume we have $k$ high-quality sensors to choose from with corresponding $\sigma_{i}$ values ranging from $\sigma_{\text{min}}$ to $\sigma_{\text{max}}$. We constrain the multimodal sensor selection problem to only consider sensor locations along the path $p$. The multimodal sensor selection problem can then be stated as:
@@ 5.4 Multi-agent IPP
The estimation of $x$ is now based on the collective measurements from all agents. The maximum a posteriori estimate from all agents measurements is:
@@ 6 Results
We provide empirical evidence of the advantages of our approach over a variety of other baselines and methods from the literature. Namely, we directly compare our method with an implementation of the mixed integer formulation presented by [2] as well as the belief MDP formulation of [17] [2, 17]. Our experiments focus on the interplay between the number of nodes in the graph, the allotted budget, and the allowed runtime to evaluate the influence on the information objective of interest. All of the results presented in this section were run on a computer with an M1 Max and 64 GB of RAM.
@@ 6.1 Experimental Setup
We consider environments modeled as grid-based graphs, but our approach scales to graphs with arbitrary structures. The graph is a $\sqrt{n}\times\sqrt{n}$ grid contained in an environment with constant edge lengths. Our experimental setup is similar to that presented by [2]; to allow for consistent comparisons we use a prediction set fixed at $m=20$ [2]. In all experiments, the start and end nodes are on diagonally opposite ends of the grid. The covariance structure of the environment is modeled using the squared exponential kernel
@@ 6.2 B-IPP as a Surrogate Objective for A-IPP
We first begin with the A-IPP problem and compare the exact mixed integer convex problem (MICP) presented in Equation 16 with the exact MICP using the B-IPP $\mathbf{tr}(\Sigma^{-1})$ objective along with the mixed integer program (MIP) formulation from [2] [2]. For the exact MICP using the B-IPP objective we are optimizing the path with respect to the B-IPP objective and then evaluating the path with respect to the A-IPP objective. Note that the MIP formulation presented by [2] only considers the A-IPP problem and does not provide an equivalent formulation for the D-IPP objective. We first compare our results on small graphs ranging from $n=4$ to $n=3844$. Our results show that on small graph sizes ($n<500$), all three of the methods perform similarly in both runtime and solution quality as shown in Figure 1.
However, when we scale up to larger graph sizes, we can see that the $\mathbf{tr}(\Sigma^{-1})$ objective can scale much more efficiently. Note that all methods are restricted to the same $120$ second time constraint and the best solution found is then returned. The MIP approach performs well for $n<500$ but fails to scale to larger graphs. Most notably, the B-IPP objective proves to be an effective surrogate for the A-IPP problem. The reason for the improved performance is twofold. First, the $\mathbf{tr}(\Sigma^{-1})$ objective is an affine function of the decision variable $z$ resulting in a mixed integer linear program. Second, by maximizing $\mathbf{tr}(\Sigma^{-1})$, we are trying to increase the magnitude of the eigenvalues of $\Sigma_{z}^{-1}$. We know that the eigenvalues of $\Sigma_{z}^{-1}$ are the reciprocals of the eigenvalues of $\Sigma_{z}$ (assuming $\Sigma_{z}$ is invertible) so increasing the eigenvalues of $\Sigma_{z}^{-1}$ will decrease the magnitude of each eigenvalue of $\Sigma_{z}$. However, increasing the sum of the eigenvalues of $\Sigma_{z}^{-1}$ will not necessarily decrease the sum of the eigenvalues of $\Sigma_{z}$ because while increasing a particular eigenvalue of $\Sigma_{z}^{-1}$ will decrease the corresponding eigenvalue of $\Sigma_{z}$, the impact on the overall trace depends on the specific eigenvalue distributions of both matrices. That is, the A and B-IPP objectives are not equivalent, but they are closely related and a meaningful performance improvement can be achieved using the B-IPP objective as a surrogate on larger problem instances.
The goal of these experiments was to isolate the comparison between these three approaches to the A-IPP objective. We now proceed to scale our experiments to even larger graphs and include both A and D-IPP objectives with additional baselines to compare against.

### v2

@@ Abstract
We consider the problem of finding an informative path through a graph, given initial and terminal nodes and a given maximum path length. We assume that a linear noise corrupted measurement is taken at each node of an underlying unknown vector that we wish to estimate. The informativeness is measured by the reduction in uncertainty in our estimate, evaluated using several Gaussian process-based metrics. We present a convex relaxation for this informative path planning problem, which we can readily solve to obtain a bound on the possible performance. We develop an approximate sequential method where the path is constructed segment by segment through dynamic programming. This involves solving an orienteering problem, with the node reward acting as a surrogate for informativeness, taking the first step, and then repeating the process. The method scales to very large problem instances and achieves performance close to the bound produced by the convex relaxation. We also demonstrate our method’s ability to handle adaptive objectives, multimodal sensing, and multi-agent variations of the informative path planning problem.
keywords:
Informative path planning , Gaussian processes , Sequential optimization
journal: Robotics and Autonomous Systems
[…]
figurec
2D
two-dimensional
3D
three-dimensional
AHRS
attitude and heading reference system
AUV
autonomous underwater vehicle
COMA
counterfactual multi-agent policy gradients
CPP
Chinese Postman Problem
DoF
degree of freedom
DVL
Doppler velocity log
FSM
finite state machine
IMU
inertial measurement unit
LBL
Long Baseline
MCM
mine countermeasures
MDP
Markov decision process
MCTS
Monte Carlo Tree Search
POMDP
partially observable Markov decision process
PRM
Probabilistic Roadmap
RL
reinforcement learning
ROI
region of interest
ROS
Robot Operating System
ROV
remotely operated vehicle
SLAM
simultaneous localization and mapping
SSE
sum of squared errors
STOMP
Stochastic Trajectory Optimization for Motion Planning
TRN
Terrain-Relative Navigation
UAV
unmanned aerial vehicle
UGV
unmanned ground vehicle
USBL
Ultra-Short Baseline
IPP
informative path planning
AIPP
adaptive informative path planning
FoV
field of view
CDF
cumulative distribution function
ML
maximum likelihood
RMSE
Root Mean Squared Error
MLL
Mean Log Loss
GP
Gaussian Process
KF
Kalman Filter
IP
Interior Point
BO
Bayesian Optimization
SE
squared exponential
UI
uncertain input
MCL
Monte Carlo Localisation
AMCL
Adaptive Monte Carlo Localisation
SSIM
structural similarity index measure
PSNR
peak signal-to-noise ratio
LPIPS
learned perceptual image patch similarity
MAE
Mean Absolute Error
RMSE
root mean squared error
AUSE
Area Under the Sparsification Error curve
mIoU
mean Intersection-over-Union
CMA-ES
covariance matrix adaptation evolution strategy
RRT
rapidly exploring random tree
RIG
rapidly exploring information gathering
AL
active learning
MLP
multi-layer perceptron
LSTM
long short-term memory
PPO
proximal policy optimization
CNN
convolutional neural network
IL
imitation learning
GPS
Global Positioning System
[…]
@@ 1 Introduction
A variety of real-world problems can be formulated as IPP problems, such as planetary rover exploration, search and rescue, and environmental monitoring. IPP merges the fields of robotics, artificial intelligence, and spatial data analysis. It carries substantial implications for a myriad of applications, where efficient data collection from a potentially hazardous or inaccessible environment is essential. However, the problem presents significant challenges. The IPP problem is known to be NP-hard due to its combinatorial nature with a vast search space for potential solutions [3]. Successful solutions typically require reasoning about the entire trajectory; myopic approaches often result in inefficient paths, causing the agent to repeat exploration of the same areas without using the full resource budget [4, 5, 6, 7, 8].
@@
The IPP problem is closely related to the optimal design of experiments and sensor selection. Joshi and Boyd presented a convex relaxation of the sensor selection problem that involves choosing a set of $k$ sensor measurements, from a set of $m$ possible or potential sensor measurements, that minimizes the error in estimating some parameters [1]. The sensor selection problem is closely related to the IPP problem. However, sensor selection does not consider the movement constraints of the agent so the rounding approach presented by Joshi and Boyd is not directly applicable to the IPP problem. Krause et al. also considered the sensor placement problem and presented several approximate solution methods [9]. Singh et al. introduced an approximate algorithm for both the single and multi-agent IPP problems using Gaussian processes as the underlying environment model, but they do not consider the adaptive IPP problem [10]. Strawser and Williams used hybrid search to solve risk-bound goal-directed planning problems, focusing on complex 3D environments and the modeling of trajectory risk [11]. Bäckström et al. focused on approximability for cost-optimal planning [12].
@@
Dutta et al. introduced a mixed-integer formulation for the discrete IPP problem in random fields and aimed to optimize over the collected measurement subset and all linear estimators [2]. The approach presented by Dutta et al. can produce globally optimal solutions on relatively small graphs; however, the mixed integer program fails to scale to larger graph sizes which is often necessary in practice for planning at higher resolutions [2]. Boström-Rost et al. formulated the problem as a mixed integer semidefinite program and used a receding horizon optimal control approach based on the information filter to produce a solution [13]. Song et al. used a receding horizon control strategy with a discrete set of motion primitives to evaluate the objective and then select the best candidate, execute the action, update the Gaussian process belief, and then replan [14].
In addition to the discrete path planning methods discussed, continuous informative path planning approaches have also been explored. These methods focus on optimizing continuous paths rather than discrete sensing locations, which is crucial for certain real-world applications where smooth trajectories and compliance with motion constraints are required. Hitz et al. used an evolutionary strategy to optimize a parameterized path in continuous space [15]. Francis et al. presented a constrained Bayesian optimization approach to plan continuous paths that inherently satisfy motion and safety constraints while balancing the reward and risk associated with each path [16]. Asgharivaskasi et al. proposed a differentiable approximation of the Shannon mutual information between a grid map and ray-based observations, enabling gradient ascent optimization in the continuous space of SE(3) sensor poses [17].
@@
Another common approach is to formulate the IPP problem as a partially observable Markov decision process (POMDP). Marchant et al. formulated the IPP problem as a POMDP and solved it using sequential Bayesian optimization through Monte Carlo tree search with upper confidence bound for trees (MCTS-UCT) [18]. This work has been extended to modify the reward function for achieving monitoring behavior that exploits areas with high gradients, and for additional reasoning over continuous action spaces through Bayesian optimization [4, 19]. Fernández et al., on the other hand, used partially observable Monte Carlo planning (POMCP) with Gaussian process beliefs for estimating quantiles of the underlying world state [20]. Their methodology proposes sample locations for a team of scientists to investigate following their exploration completion. Rückin et al. presented a hybrid approach that combines tree search with an offline-learned neural network to predict informative sensing actions [21]. Ott et al. focused on the multimodal sensing extension of the IPP problem where the agent has multiple sensors to choose from and solved it with tree search [22].
@@
These approaches all produce reasonable solutions on graphs with a relatively small number of nodes; however, they struggle to scale to large graphs with higher resolution which is necessary for many real-world applications. Additionally, MIPs are not well-suited to handle adaptive objectives where the objective function depends on the history of measurement values received. Our results demonstrate that our approach is able to scale to larger problem instances for the IPP problem, while also handling adaptive objectives such as the expected improvement metric, and accounting for multimodal sensing and multi-agent variations of the IPP problem [23, 22]. We directly compare our approach with those of Dutta et al. and Ott et al. [2, 22].
@@ 2.1 Feasible Paths
We model the environment as a directed graph with vertices $V=\{1,\dots,n\}$ and edges $\mathcal{E}\subseteq V\times V$. Our goal is to find a feasible path $p$ from the start vertex $s$ to the goal vertex $g$ where we assume that such a feasible path exists. We denote a path of length $L$ as $p=(i_{1},i_{2},\dots,i_{L})$ where $i_{1}=s$ and $i_{L}=g$. A feasible path is one that starts at $s$ and ends at $g$ while remaining within the given maximum path length $L^{\text{max}}$; that is, $L\leq L^{\text{max}}$.
@@ 2.2 Sensing Along a Path
$\hat{x}=\bar{x}+\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\sigma^{-2}\sum_{k=1}^{L}(y_{k}-\bar{y}_{k})a_{i_{k}}$ (3)
@@
$\Sigma=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}$ (4)
where $\bar{x}$ is the prior mean and $\bar{y}_{k}=a_{i_{k}}^{T}\bar{x}$. Note that this is exactly the Gaussian process mean and covariance as shown in D.
@@ 2.3 Measures of Informativeness
Note that each scalar measure of informativeness is a function of the eigenvalues of $\Sigma$. We will denote the measure of informativeness of a path as $\phi(p)$ where smaller values of $\phi(p)$ correspond to more informative paths. Other measures of informativeness such as adaptive objectives will be discussed in section 5.2 and we also provide a formulation of the mutual information in E.
@@ 3 Mixed Integer Convex Formulation
$\Sigma_{z}=\left(\sigma^{-2}\sum_{i=1}^{n}\left(\sum_{j\in N_{out}(i)}z_{ij}\right)a_{i}a_{i}^{T}+\Sigma_{x}^{-1}\right)^{-1}$ (15)
and can write our parameterized objective as $\phi(z)$. We then have the following convex mixed integer problem:
@@
While it is not immediately obvious, the $\phi(z)$ objective is convex in both the A, B, and D-IPP objectives introduced in section 2.3 as discussed in section 4.6.2 of Boyd and Vandenberghe. The problem would be entirely convex if it were not for the integer constraints on $z$ in equation 14.
@@ 5.2 Additional Objectives
$\mathbb{E}[I_{p}]_{j}=(y_{\text{min}}-\hat{x}_{j})\Phi\left(\frac{y_{\text{min}}-\hat{x}_{j}}{\hat{\sigma}_{j}}\right)+\hat{\sigma}_{j}^{2}\mathcal{N}(y_{\text{min}}\mid\hat{x}_{j},\hat{\sigma}_{j}^{2})$ (22)
where $y_{\text{min}}$ is the current minimum value found so far, and $\Phi(\cdot)$ is the cumulative distribution function [24]. Our problem then becomes
minimize $\sum_{j=1}^{m}\mathbb{E}[I_{p}]_{j}$
@@
This is a non-convex objective and is not easily handled by current optimizers. However, our approximate sequential path optimization approach can handle this objective by computing $r_{j}=\mathbb{E}[I_{p}]_{j}$ for each of the $m$ prediction vertices in the environment with the current values of $\hat{x}$ and $\Sigma$ obtained from the current path $p_{t}$. Note that we want to visit locations in the environment with high expected improvement to minimize the net expected improvement left in the environment so there is no negative sign in the expression for $r_{j}$
@@ 5.3 Multimodal Sensing
To solve the AIPPMS problem, we first solve the AIPP problem with our approximate sequential path optimization approach and the respective objective presented in section 4. We then use the resulting path as the starting point for the multimodal sensor selection problem. The multimodal sensor selection problem is a variant of the sensor selection problem presented by Joshi and Boyd. In this case, we assume we have $k$ high-quality sensors to choose from with corresponding $\sigma_{i}$ values ranging from $\sigma_{\text{min}}$ to $\sigma_{\text{max}}$. We constrain the multimodal sensor selection problem to only consider sensor locations along the path $p$. The multimodal sensor selection problem can then be stated as:
@@ 5.4 Multi-agent IPP
The estimation of $x$ is now based on the collective measurements from all agents similar to the sequential allocation method proposed by Singh et al. for multi-agent exploration [10]. The maximum a posteriori estimate from all agents measurements is:
@@ 6 Results
We provide empirical evidence of the advantages of our approach over a variety of other baselines and methods from the literature. Namely, we directly compare our method with an implementation of the mixed integer formulation presented by Dutta et al. as well as the belief MDP formulation of Ott et al. [2, 22]. Our experiments focus on the interplay between the number of nodes in the graph, the allotted budget, and the allowed runtime to evaluate the influence on the information objective of interest. All of the results presented in this section were run on a computer with an M1 Max and 64 GB of RAM.
@@ 6.1 Experimental Setup
We consider environments modeled as grid-based graphs, but our approach scales to graphs with arbitrary structures. The graph is a $\sqrt{n}\times\sqrt{n}$ grid contained in an environment with constant edge lengths. Our experimental setup is similar to that presented by Dutta et al.; to allow for consistent comparisons we use a prediction set fixed at $m=20$ [2]. In all experiments, the start and end nodes are on diagonally opposite ends of the grid. The covariance structure of the environment is modeled using the squared exponential kernel
@@ 6.2 B-IPP as a Surrogate Objective for A-IPP
We first begin with the A-IPP problem and compare the exact mixed integer convex problem (MICP) presented in equation 16 with the exact MICP using the B-IPP $\mathbf{tr}(\Sigma^{-1})$ objective along with the mixed integer program (MIP) formulation from Dutta et al. [2]. For the exact MICP using the B-IPP objective we are optimizing the path with respect to the B-IPP objective and then evaluating the path with respect to the A-IPP objective. Note that the MIP formulation presented by Dutta et al. only considers the A-IPP problem and does not provide an equivalent formulation for the D-IPP objective. We first compare our results on small graphs ranging from $n=4$ to $n=3844$. Our results show that on small graph sizes ($n<500$), all three of the methods perform similarly in both runtime and solution quality as shown in figure 1.
However, when we scale up to larger graph sizes, we can see that the $\mathbf{tr}(\Sigma^{-1})$ objective can scale much more efficiently. Note that all methods are restricted to the same $120$ second time constraint and the best solution found is then returned. The MIP approach performs well for $n<500$ but fails to scale to larger graphs. Most notably, the B-IPP objective proves to be an effective surrogate for the A-IPP problem. The reason for the improved performance is twofold. First, the $\mathbf{tr}(\Sigma^{-1})$ objective is an affine function of the decision variable $z$ resulting in a mixed integer linear program. Second, by maximizing $\mathbf{tr}(\Sigma^{-1})$, we are trying to increase the magnitude of the eigenvalues of $\Sigma_{z}^{-1}$. We know that the eigenvalues of $\Sigma_{z}^{-1}$ are the reciprocals of the eigenvalues of $\Sigma_{z}$ (assuming $\Sigma_{z}$ is invertible). Therefore, increasing the eigenvalues of $\Sigma_{z}^{-1}$ will decrease the corresponding eigenvalues of $\Sigma_{z}$. While maximizing $\mathbf{tr}(\Sigma_{z}^{-1})$ indirectly influences the eigenvalues of $\Sigma_{z}$, the trace of $\Sigma_{z}$ and the trace of $\Sigma_{z}^{-1}$ are not simply reciprocals of each other. That is, the A and B-IPP objectives are not equivalent, but they are closely related and a meaningful performance improvement can be achieved using the B-IPP objective as a surrogate on larger problem instances. The goal of these experiments was to isolate the comparison between these three approaches to the A-IPP objective. We now proceed to scale our experiments to even larger graphs and include both A and D-IPP objectives with additional baselines to compare against.